OpenStudy (anonymous):

0-theta ...if sin0 +sin^2 0 = 1 ...then proove cos^2 0+cos^4 0 =1

OpenStudy (anonymous):

So you mean if \[\sin\theta+\sin^2\theta=1\] show that \[\cos^2\theta+\cos^4\theta=1~~?\] The first equation can be solved for \(\theta\): \[\begin{align*}\sin\theta+\sin^2\theta&=1\\ \sin^2\theta+\sin\theta-1&=0\\ \sin^2\theta+\sin\theta+\frac{1}{4}-\frac{5}{4}&=0\\ \left(\sin\theta+\frac{1}{2}\right)^2&=\frac{5}{4}\\ \sin\theta+\frac{1}{2}&=\pm\frac{\sqrt5}{2}\\ \sin\theta&=\frac{-1\pm\sqrt5}{2} \end{align*}\] In fact, you don't have to solve directly for \(\theta\) to continue. \(\sin\theta\) will do fine. Recall the identity: \(\sin^2\theta+\cos^2\theta=1\). The second equation is then \[\begin{align*}\cos^2\theta+\cos^4\theta&=\left(1-\sin^2\theta\right)+\left(1-\sin^2\theta\right)^2 \end{align*}\] You have two values for \(\sin\theta\) that give \[\sin^2\theta=\frac{3-\sqrt5}{2},~\frac{3+\sqrt5}{2}\] Plugging into the equation above, you have \[\begin{align*}\cos^2\theta+\cos^4\theta&=\left(1-\frac{3-\sqrt5}{2}\right)+\left(1-\frac{3-\sqrt5}{2}\right)^2\\ &\text{and}\\ \cos^2\theta+\cos^4\theta&=\left(1-\frac{3+\sqrt5}{2}\right)+\left(1-\frac{3+\sqrt5}{2}\right)^2 \end{align*}\] Verify that both right hand sides reduce to 1.